OpenStudy (anonymous):

(3x + 4)(3x + 7) = 0

jimthompson5910 (jim_thompson5910):

if (3x + 4)(3x + 7) = 0 then 3x+4 = 0 or 3x+7 = 0

jimthompson5910 (jim_thompson5910):

solve each equation separately for x to get your two answers